# Computing Wasserstein Barycenters through Gradient Flows

### v2

@@ 2.4 Gradient Flows
In the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional. A gradient flow is the solution to the differential equation,
$\begin{cases}\dot{x}(t)=-\nabla F(x(t))&\text{for }t>0,\\ x(0)=x_{0},\end{cases}$ (11)
where $x_{0}$ is the initial condition. In $\mathcal{P}_{p,\text{ac}}(\Omega)$, gradient flows correspond to the continuity equation,
$\partial_{t}P_{t}=-\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})),$ (12)
where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures. Usually, these functionals take the following form [santambrogio2015optimal],
$\mathbb{F}(P)=$ $\underbrace{\int G(P(z))dz}_{\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\mathbb{V}(P)}+$ (13)
$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\mathbb{U}(P)},$
where $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ are the internal, potential, and interaction energies, respectively. Here, $G:\mathbb{R}\rightarrow\mathbb{R}$ is convex and superlinear, and $V:\Omega\rightarrow\mathbb{R}$ and $U:\Omega^{2}\rightarrow\mathbb{R}$ are convex and sufficiently smooth. Henceforth, we denote $\mathbb{F}^{\star}=\text{inf}_{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}\mathbb{F}(P)$.
@@ 3 Wasserstein Barycenters as Gradient Flows
In this section, we describe a new method for computing empirical and Gaussian mixture Wasserstein barycenters. Our main idea is using the functional:
$\mathbb{F}(P)=\mathbb{B}_{\mathcal{Q}}(P)+\mathbb{G}(P)+\mathbb{V}(P)+\mathbb{U}(P),$ (14)
where $\mathbb{B}_{\mathcal{Q}}(P)$ is the barycenter objective defined in equation 7, and $\mathbb{G},\mathbb{V}$ and $\mathbb{U}$ are the energies defined in equation 13.
3.1 Empirical Flow
In the empirical case, the barycenter is approximated through a finite set of particles (see equation 1) $\{z_{i}^{(P)}\}_{i=1}^{n}$. Hence, we may rewrite equation 14 as,
$\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{min}}\,\underbrace{\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}d(z_{i}^{(P)},z_{j}^{(Q_{k})})^{p}}_{\mathbb{B}_{\mathcal{Q}}(\hat{P})}$+
$\underbrace{\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)})}_{\mathbb{V}(\hat{P})}+\underbrace{\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)})}_{\mathbb{U}(\hat{P})}$ (15)
Here, we omit the internal energy term, $\mathbb{G}(P)$, as it is not defined for empirical measures. We solve the previous optimization problem through a block-coordinate descent strategy, that is,
$\gamma_{k}$ $=\underset{\gamma\in\Gamma(\hat{P}_{\tau},\hat{Q}_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{i,j}d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{p},$
$z_{\tau+1,i}^{(P)}$ $=z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}).$ (16)
In equation 16, $\gamma_{k}$ is used to compute $\mathbb{B}_{\mathcal{Q}}$ and is updated for each $\tau$. We provide in Algorithm 1 the pseudo-code for this strategy. Our algorithm assumes access to $Q_{k}$ through sampling, that is, we do not necessarily have access to all samples of $Q_{k}$ at once.
Algorithm 1 Empirical barycenter Wasserstein flow using Gradient Descent.
@@ 3.1 Empirical Flow
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
5:  for $k\leftarrow 1$ to $K$ do
6:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i.i.d. from $Q_{k}$.
7:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{p}(\hat{P}_{\tau},\hat{Q}_{k})^{p}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\tau})$
10:end for
Remark 3.1.
In line 5 of Algorithm 1, $\hat{Q}_{k}$ denotes the empirical measure (c.f., equation 1) supported on the mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ sampled from $Q_{k}$.
3.2 Gaussian Mixture Flow
In the case of GMMs, we flow the parameters $\theta=\{(\pi_{k},\mu_{k},\Sigma_{k})\}_{k=1}^{K}$ of $P_{\theta}=\sum\pi_{i}\mathcal{N}(\cdot|\mu_{i},\Sigma_{i})$. Similarly to equation 16, we have:
$\omega_{k}$ $=\underset{\omega\in\Gamma(\pi,\pi_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\omega_{i,j}\mathbb{W}_{2}(P_{\theta_{\tau},i},Q_{k,j})^{2},$
$\theta_{\tau+1,i}$ $=\theta_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}}).$ (17)
Here, $\omega_{k}$ is a transport plan between GMM components. In general, the energy functionals do not have closed form with respect to the GMM parameters. However, we can estimate them using Monte-Carlo and the reparametrization trick, for instance,
$z_{i}$ $=L_{i}\varepsilon+\mu_{i},\quad i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id}),$ (18)
This strategy allows us to propagate the gradients to the parameters of $P_{\tau}$. We show the overall approach in Algorithm 2.
Algorithm 2 GMM barycenter flow using Gradient Descent.
@@ 3.2 Gaussian Mixture Flow
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  Sample $\{z_{i}=L_{i}\varepsilon+\mu_{i}:i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id})\}$,
5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{G}(\hat{P}_{\tau})+\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
@@
7:   $\mathbb{F}(P_{\theta_{\tau}})\leftarrow\mathbb{F}(P_{\theta_{\tau}})+\lambda_{k}\mathbb{MW}_{2}(P_{\theta_{\tau}},Q_{k})^{2}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}})$
10:end for
Covariance matrix parametrization. Optimizing $\mathbb{F}$ with respect to $\Sigma_{k}$ is challenging, due to the symmetric positive definite constraint. We enforce this constraint by writing $\Sigma_{k}=L_{k}L_{k}^{T}$, where $L_{k}$ is the lower-triangular Cholesky factor. There are two advantages to this strategy. First, optimizing with respect to $L_{k}$ is empirically more stable than $\Sigma_{k}$. Second, we can use $L_{k}$ directly in the reparametrization trick in equation 18.
@@ 3.3 Flows over Joint Measures
In [cuturi2014fast], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work. As we show in our experiments, in machine learning applications one encounters $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Whenever we deal with labeled measures (i.e., $z=(x,y)$), we adopt,
@@
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in [montesuma2021wasserstein] and [alvarez2021dataset], is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
@@
thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$. From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
For GMMs, we equip their components with labels $\nu_{k}\in\Delta_{n_{\text{classes}}}$. Using the ground-cost in equation 27, we are able to write the $\mathbb{MW}_{2}(P,Q)^{2}$ as,
$\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}^{\star}(\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\beta\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}^{2}),$
where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal. This approach was first proposed in [montesuma2024lighter], but the authors relied on heuristic arguments to justify this modeling choice. We present the following proposition justifying the label term,
Proposition 3.1.
Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be two GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$. Let the ground cost $c$ be,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
Functionals for joint measures. One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by [alvarez2021dataset] for transfer learning problems. Here we propose the following functionals,
$V_{E}(z)$ $=-\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (20)
$U_{R}(z,z^{\prime})$ $=\begin{cases}h(d(x,x^{\prime}))&\text{if }y\neq y^{\prime},\\ 0&\text{otherwise}\end{cases},$ (21)
where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below. For instance, in our experiments we use the hinge loss, $h(u)=\text{max}(0,\text{margin}-d)$, where $\text{margin}\geq 0$ is a margin parameter. Equations 20 and 21 correspond to entropy and repulsion terms. The first functional penalizes barycenters that have fuzzy labels. The second functional encourages classes to be well separated.

### v3

@@ 2.3 Gradient Flows
Our method relies on the gradient flow interpretation of the minimization in equation 8. Starting from the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional. A gradient flow is the solution to the Ordinary Differential Equation (ODE),
$\dot{x}(t)=-\nabla F(x(t))\text{, subject to }x(0)=x_{0}.$ (9)
This framework extends to probability measures. Let $\mathbb{F}:\mathcal{P}_{2}(\Omega)\rightarrow\mathbb{R}$ be a functional over $\mathcal{P}_{2}(\Omega)$. The evolution of a curve $\{P_{t}\}_{t\geq 0}$ in $\mathcal{P}_{2}(\Omega)$ is given by the continuity equation
$\partial_{t}P_{t}=-\text{div}(P_{t}v_{t}),$ (10)
where $v_{t}:\Omega\rightarrow\Omega$ is a velocity field, and div is the divergence operator, understood in a measure theoretic sense. Choosing $v_{t}=-\nabla_{W}\mathbb{F}(P_{t})$ yields the gradient flow of $\mathbb{F}$, that is, the measure theoretic analogue of equation 9. Here, $\nabla_{W}$ denotes the Wasserstein gradient of the functional $\mathbb{F}$ (c.f. [chewi2024statistical, Section 5.4] and appendix C.1).
The structure of $\mathbb{F}$ determines the dynamics of $P_{t}$. Following the classical OT theory [santambrogio2015optimal, Chapter 7], we consider the functional $\mathbb{F}(P)=\mathbb{B}(P)+\mathbb{R}(P)$, where,
$\mathbb{R}(P)=$ $\underbrace{\int G(p(z))dz}_{\text{Internal energy }\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\text{Potential energy }\mathbb{V}(P)}+$ (11)
$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\text{Interaction energy }\mathbb{U}(P)}$
where $p:\Omega\rightarrow\mathbb{R}$ denotes the density of $P$, $G:\mathbb{R}\rightarrow\mathbb{R}$, $V:\Omega\rightarrow\mathbb{R}$, and $U:\Omega\times\Omega\rightarrow\mathbb{R}$. For example, when $(s)=s\log s, 𝔾\mathbb{G}$ is the entropy functional of [chizat2025doubly], associated with diffusion dynamics. $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ introduce diffusion, drift, and pairwise behaviors in the flow, respectively. See Section 3.3 for further examples.
@@ 3 Wasserstein Barycenters as Gradient Flows
In this section, we describe a new method for computing empirical OT barycenters. We use,
[…]
First, we derive the empirical version of the continuity equation (c.f., Equation 10). For $\hat{P}_{t}\in\text{Emp}_{n}(\Omega)$:
$\partial_{t}\hat{P}_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\partial_{t}\delta_{z_{t,i}^{(P)}}=-\dfrac{1}{n}\sum_{i=1}^{n}\text{div}(\dot{z}_{t,i}^{(P)}\delta_{z_{t,i}^{(P)}}),$ (13)
$=-\text{div}(\hat{P}_{t}v_{t}),$
where we identified $v_{t}(z_{t,i}^{(P)})=\dot{z}_{t,i}^{(P)}$. Second, we discretize the gradient flow in time, through an index $\tau=\alpha t$, $\alpha>0$. Using the forward Euler scheme,
$\dot{z}_{t,i}^{(P)}=\frac{z_{\tau+1,i}^{(P)}-z_{\tau,i}^{(P)}}{\alpha}\rightarrow z_{\tau+1,i}^{(P)}=z_{\tau,i}^{(P)}+\alpha v_{\tau,i},$ (14)
where $v_{\tau,i}=v_{\tau}(z_{\tau,i}^{(P)})$. We can obtain different update strategies based on the velocity field,
$v_{\tau,i}$ $=-\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau}),$ (15)
$v_{\tau,i}$ $=-\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau})+\sqrt{2\alpha\eta_{G}}\eta_{\tau,i},$ (16)
$v_{\tau,i}$ $=u_{\tau,i},\quad u_{\tau+1,i}=\beta u_{\tau,i}-\alpha\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau}).$ (17)
Equations 15, 16 and 17 correspond to the steepest descent, Langevin dynamics, and momentum descent strategies, respectively. $\eta_{G}\geq 0$ denotes the diffusion coefficient. The Langevin dynamics version of the flow corresponds to the NPGD dynamics of chizat2025doubly. Furthermore, for empirical measures, the functionals in Equation 12 are tractable, and take the following form,
[…]
where $\hat{p}$ denotes the density of $\hat{P}$. Since $\hat{P}$ is empirical it is, in general, difficult to compute $\mathbb{G}$. A notable exception is the entropy functional (i.e., $G(s)=s\log s$), which adds a diffusion term to the gradient flow dynamics. This corresponds to Langevin dynamics [chizat2025doubly] (Equation 16). For general choices of $G$, one needs to estimate the density $\hat{p}$ at each iteration of the empirical flow.
Algorithm 1 Empirical Time Discretized Wasserstein Gradient Flow
@@
4:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})+\mathbb{G}(\hat{P}_{\tau})$
@@
7:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i .i.d. from $Q_{k}$.
8:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{2,\epsilon}(\hat{P}_{\tau},\hat{Q}_{k})^{2}$
9:  end for
10:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}+\alpha v_{\tau,i}$
11:end for
An immediate consequence of Equation 18 is that our gradient flow formulation encompasses several previous works. For instance, with $G=U=V=0$, we have the classical free-support barycenters of cuturi2014fast. Furthermore, for the entropy functional and $\epsilon>0$, we have the NPGD algorithm of chizat2025doubly. Finally, as we show in Section 3.2 below, with a specific metric on $\Omega=\mathcal{X}\times\mathcal{Y}$, we retrieve the algorithm of montesuma2023multi.
3.1 Scaling with Mini-Batch OT
The computation of $\mathbb{B}_{\epsilon}$ in the discrete approach [cuturi2014fast] involves an OT problem between the complete set of samples in the input measures. Therefore, when the input measures are large scale, the barycenter problem becomes infeasible. We circumvent this issue with mini-batch OT [fatras2021minibatch].
The main idea behind our scalable barycenter algorithm is accessing $\hat{Q}_{k}\in\mathcal{Q}$ through sampling. At each iteration of Algorithm 1, we sample a mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ of $m$ samples from each input measure (cf. line 5 in Algorithm 1). Since we sample the same number of samples from each measure, we can actually vectorize the computation of the $K$ OT problems in the evaluation of $\mathbb{W}_{2,\epsilon}$.
Let $\{z_{\tau,i}^{(P)}\}_{i=1}^{n}$ be the barycenter support, and $\{z_{j}^{(Q_{k})}\}_{j=1}^{m}$ be a mini-batch from $Q_{k}$. The Gibbs kernel $E_{k,i,j}=e^{-(C_{k,i,j}/\epsilon)}$, and $C_{k,i,j}=d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{2}$ are arrays with shape $(K,n,m)$. The Sinkhorn iterations become,
$\psi_{k}^{(\ell+1)}=\dfrac{b}{E_{k}^{T}\phi_{k}^{(\ell)}}\text{, and }\phi_{k}^{(\ell+1)}=\dfrac{a_{k}}{E_{k}\psi_{k}^{(\ell+1)}},$ (19)
where the division operation should be understood coordinate-wise. Furthermore, $b_{j}=n^{-1}$ and $a_{k,i}=m^{-1}$ is the importance of each sample in the barycenter, and input measure support, respectively. For simplicity, we assume uniform weights. For $L\in\mathbb{N}$ iterations, we can retrieve $\gamma_{\epsilon,k}^{\star}=\text{diag}(\phi_{k}^{(L)})E_{k}\text{diag}(\psi_{k}^{(L)})$. Since each $\gamma_{\epsilon,k}^{\star}$ has shape $(n,m)$, we can actually stack it in an array of shape $(K,n,m)$, similarly to $E$ and $C$.
The main advantage of equation 19 is that the $K$ OT problems can be vectorized. This is possible because we draw the same number of samples from each input measure, making it possible to vectorize the Sinkhorn iterations. For instance, a similar strategy is possible for measures over a fixed grid [peyre2019computational, Remark 4.16]. The OT plan $\gamma_{\epsilon}$, obtained at the end of the Sinkhorn iterations, is used to compute $\mathbb{B}_{\epsilon}$, in lines 4 through 7 in Algorithm 1.
@@ 3.2 Flows over Joint Measures
An interesting advantage of the gradient flow formalism is that the only necessary ingredient is a differentiable metric over $\Omega$. Therefore, we generalize the setting of cuturi2014fast, who assumed $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. As we show in our experiments, in machine learning applications one has $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Let $z=(x,y),x\in\mathcal{X},y\in\mathcal{Y}$,
@@
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in montesuma2021wasserstein and alvarez2021dataset, is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{n_{\text{classes}}}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
@@
thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$, $x\in\mathbb{R}^{d}$, $\ell\in\mathbb{R}^{n_{\text{classes}}}$. From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{\text{hard},i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
3.3 Task-Aware Regularizing Functionals
One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by alvarez2021dataset for transfer learning problems. Here we propose the following functionals,
$V_{E}(z)$ $=-\eta_{V}\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (21)
$U_{R}(z,z^{\prime})$ $=\begin{cases}\eta_{U}h(d(x,x^{\prime}))&\text{if }y_{\text{hard}}\neq y_{\text{hard}}^{\prime},\\ 0&\text{otherwise}\end{cases},$ (22)
where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below, and $\eta_{V}>0$ and $\eta_{U}>0$ are scaling parameters. For instance, in our experiments we use the hinge loss, $h(u)=\text{max}(0,\text{margin}-d)$, where $\text{margin}\geq 0$ is a margin parameter. Equations 21 and 22 correspond to entropy and repulsion, respectively. The first functional penalizes barycenters that have fuzzy labels. The second functional encourages classes to be well separated.